How do you use the Systems of Linear Equation but with the Elimination Method? Please explain with the question, x-10y=60 , x+-14y=12. Thankyou!
Answer by ankor@dixie-net.com(22740)

x - 10y = 60
x - 14y = 12
------------------Subtraction eliminates x, find y
4y = 48
y = 48/4
y = 12
:
Use the first equation to find x, replace y with 12
x - 10(12) = 60
x - 120 = 60
x = 60 + 120
x = 180
:
Check the solutions in the 2nd equation
x - 14y = 12
replace x and y
180 - 14(12) = 12
180 - 168 = 12